The Federal Reserve System publishes data on family income based on its Survey of Consumer Finances. When the head of the household has a college degree, the mean before-tax family income is $ 84,800. Suppose that 64% of the before-tax family incomes when the head of the household has a college degree are between $75,000 and $94,600 and that these incomes are normally distributed. What is the standard deviation of before-tax family incomes when the head of the household has a college degree?

(Round the value of z to 2 decimal places. Round your answer to 2 decimal places.)
Answer #1

P(75000 < X < 94600) = 0.64

75000 - or, PC 21 X-u 94600 M ) = 0.64 <-

75000 - 84800 or, PG 94600 - 84800 = 0.64

or, P=9800 <z <3800) = 0.64

9800 or, P(Z < P(Z <- -9800 -) = 0.64

or, 2 * P(Z < 9800 -)-1=0.64

or, P(Z < 9800 -) = 0.82

9800 OTO -= 0.92

9800 G = 10652.17 10.92 (ans)